OpenStudy (anonymous):

Last month , you made long-distance calls to two of your friends . One call lasted 8 minutes longer than the other. According to the phone bill, the two call lasted a total of 42 min. How long was each call?

OpenStudy (anonymous):

we must set the total equal to the component parts in an equation. Total = (call1) + (call2) 42min = (x) + (x+8) where x = the shorter call. Now, let's simplify. 42 = 2x + 8 34 = 2x 17 = x The shorter call was 17 minutes, and the longer (x + 8)=(17+8) was 25 minutes long.